Perform the operations. Simplify answers, if possible. Assume that no denominators are 0
2y/9 + y/3
/= fractions
I know the answer is 5y/9. Can you show me the steps?

Answer by crousejohn(2)

2y/9+y/3
Just like adding any fraction start by finding lowest common denominator
LCD=9
To get y/3 with 9 as it's denominator without changing it's value multiply by 3/3. 3/3=1
(y/3)(3/3)=3y/9
Now you can add them, so
2y/9+3y/9=5y/9